OpenStudy (anonymous):

Find dy/dx in terms of x and y if cos^2(6 y) + sin^2(6 y) = y + 2

zepdrix (zepdrix):

\[\Large \cos^26y+\sin^26y\quad=\quad y+2\]We want dy/dx? Hmm ok.. let's see..

zepdrix (zepdrix):

Can we cheat and use our trig identity? +_+\[\Large \color{royalblue}{\cos^2x+\sin^2x\quad=\quad 1}\]Or do they want us to work it out the long way? :)

zepdrix (zepdrix):

Understand what I mean by cheating? :D\[\Large \color{teal}{\cos^26y+\sin^26y}\quad=\quad y+2\]becomes,\[\Large \color{teal}{1}\quad=\quad y+2\]